# Coordinate Matrix Machine: A Human-level Concept Learning to Classify Very Similar Documents

### v1

@@ Abstract
Contribution: In this paper, we present the Coordinate Matrix Machine (CM2). This purpose-built small model augments human intelligence by learning document structures and using this information to classify documents. While modern “Red AI” trends rely on massive pre-training and energy-intensive GPU infrastructure, CM2 is designed as a Green AI solution. It achieves human-level concept learning by identifying only the structural “important features” a human would consider, allowing it to classify very similar documents using only one sample per class.
@@ 1 Introduction
Human-level concept learning is a relatively newer area of research. The goal is to provide a solution to a problem that is easy for humans to do, yet still challenging for machines despite their computational power. People can learn new concepts from only one or a very few samples, whereas machine learning methods require many examples to discover correlations and understand features. The reason is that people subconsciously select the important features and then generalize to other distinct areas [lake2015, mnih2015].
In this paper, we deploy a human-level concept learning approach for document classification. Document classification is a common task in machine learning and Natural Language Processing (NLP), in which a document is assigned to one or more classes. Current approaches rely heavily on the document context to classify documents [jones1972, power2010]. Identifying the relevant topic is the primary motivator for such document classification [hassan2017, kotenko2015, lin2006]. These models assume we have access to ample labelled data and that the document context is sufficiently informative to distinguish between classes. Some other methods use document image information, thereby exacerbating the labeling process. Since images are high-dimensional data, labeling so many training samples is not feasible in most real-world problems [krizhevsky2012].
@@
Our aim is to develop a coordinate matrix-based framework (Coordinate Matrix Machine, CM2) as a purpose-built small model that achieves human-level concept learning without the prohibitive costs of Large Language Models (LLMs). Unlike ’Red AI’ trends that rely on massive pre-training and energy-intensive GPU infrastructure, CM2 is designed as a Green AI solution. It prioritizes computational efficiency and environmental sustainability by mimicking human structural identification rather than processing exhaustive semantic vectors. This enables a high-precision, explainable alternative that is both economically viable and runs on standard CPU hardware.
@@ 1.1 Challenges
  • There are way too many class labels at the template level. For the five banks included in this experiment, 53 templates (classes) were available. There are not enough representative samples for each class.
  • All bank statements are very similar, as they all include many transaction descriptions, dates, and amounts.
@@
When classifying statements using the Term Frequency Vectorizer with Logistic Regression, the model with the best parameters achieved an $F$-measure of 79% on the templates.
@@ 1.2 Our Contribution
The work closest to ours in approach is that of Lake et al. [lake2015]. They have argued that in most cases, people can learn a new concept from just one or a handful of examples. At the same time, typical machine learning algorithms require hundreds of examples to perform similarly. They aim to address a different problem: recognising handwritten characters from image data. In this paper, we aim to classify structured documents from text data. Although the issues differ, both works use human-level concept learning by examining how humans solve the problem. Both approaches only use one sample per class and learn from that sample.
@@ 1.3 Advantages
    High Accuracy with Minimal Data (One-Shot Learning): The algorithm achieves human-level concept learning, successfully classifying documents using only a single sample per class. It outperforms traditional machine learning and advanced deep learning models that typically require hundreds or thousands of labeled examples to achieve comparable accuracy.
@@
    Geometric and Structural Intelligence: Unlike standard transformer-based models (like BERT or GPT) that process text as a linear sequence of tokens, CM2 utilizes a coordinate matrix to capture the physical geometry of the document. This makes it significantly more effective for formal documents (bank statements, invoices) where the spatial position of elements is more informative than semantic flow.
@@
    Optimized for CPU-Only Environments: While modern NLP trends necessitate expensive, high-end GPU clusters, CM2 is purpose-built to run on standard consumer-grade CPUs. By using static embeddings, such as GloVe, with $O(1)$ lookup latency rather than complex transformer passes, it ensures near-zero inference time on basic hardware.
@@
    Economic Viability: By eliminating the need for expensive GPU infrastructure, high-cost LLM API subscriptions, and the labor-intensive process of large-scale data labeling, $CM^{2}$ provides a significantly cheaper solution for enterprise-level deployment.
@@
    Generic, Expandable, and Extendable: The coordinate matrix framework is highly versatile. It can be easily adapted to new document templates or extended to include additional feature markers without requiring the model to be “re-trained” on a massive corpus.
@@ 2 Related Work
Once the feature vector is identified, the third step is to apply a machine learning model for classification. Naïve Bayes is one of the probability-based classifiers [kotenko2015]. Decision tree-based approaches have also been used for various purposes, such as blocking inappropriate web content [liu2017]. Support Vector Machines (SVM) are also widely used for document classification [farhoodi2010, lin2006]. There are multiple approaches under artificial/deep neural networks. Hassan and Mahmood [hassan2017] applied Convolution Neural Networks (CNN) for document classification. Unlike most machine learning models, Recurrent Neural Networks (RNNs) consider the sequence of word occurrences; therefore, documents with similar words can yield different outputs because of word order [banerjee2019, kumar2019].

### v2

@@ Abstract
Contribution: In this paper, we introduce Coordinate Matrix Machine (CM2), a structure-aware document intelligence model designed to bridge the gap between semantic content and spatial layout. Unlike traditional NLP approaches that prioritize sequential text, CM2 is particularly effective for low-data, layout-sensitive tasks, enabling high-precision performance when structural geometry is paramount. While modern “Red AI” trends rely on massive pre-training and energy-intensive GPU infrastructure, CM2 is designed as a Green AI solution. It achieves human-level concept learning by identifying only the structural “important features” a human would consider, allowing it to classify very similar documents using only one sample per class.
@@ 1 Introduction
Human-level concept learning is a relatively less explored area of research. The goal is to provide a solution to a problem that is easy for humans to do, yet still challenging for machines despite their computational power. People can learn new concepts from only one or a very few samples, whereas machine learning methods require many examples to discover correlations and understand features. The reason is that people subconsciously select the important features and then generalize to other distinct areas [lake2015, mnih2015].
In this paper, we deploy a human-level concept learning approach for document classification. Document classification is a common task in machine learning and Natural Language Processing (NLP), in which a document is assigned to one or more classes. Current approaches rely heavily on the document context to classify documents [jones1972, power2010]. Identifying the relevant topic is the primary motivator for such document classification [hassan2017, kotenko2015, lin2006]. These models assume we have access to ample labelled data and that the document context is sufficiently informative to distinguish between classes. Some other methods use document image information, thereby exacerbating the labeling process. Since images are high-dimensional data, labeling so many training samples is not feasible in most real-world problems [krizhevsky2012]. While most approaches treat text in isolation, our solution rests on the premise that the meaning of a structured document is derived from both the text and its coordinates. This dual modality allows the model to leverage spatial grounding that purely text-based models overlook.
@@
Our aim is to develop a coordinate matrix-based framework (Coordinate Matrix Machine, CM2) as a purpose-built small model that achieves human-level concept learning without the prohibitive costs of Large Language Models (LLMs). Unlike ‘Red AI’ trends that rely on massive pre-training and energy-intensive GPU infrastructure, CM2 is designed as a Green AI solution. It prioritizes computational efficiency and environmental sustainability by mimicking human structural identification rather than processing exhaustive semantic vectors. This enables a high-precision, explainable alternative that is both economically viable and runs on standard CPU hardware.
@@ 1.1 Challenges
  • There are way too many class labels at the template level. For the five Australian banks included in this experiment, 53 templates (classes) were available. There are not enough representative samples for each class.
  • All bank statements are very similar in context, as they include many transaction descriptions, dates, and amounts.
@@
When classifying statements using the Term Frequency Vectorizer with Logistic Regression, the model with the best parameters achieved an $F$-measure of 79% on the templates over 53 templates in an 80-20 training-test split.
@@ 1.2 Our Contribution
The work closest to ours in approach is that of Lake et al. [lake2015]. They have argued that in most cases, people can learn a new concept from just one or a handful of examples. At the same time, typical machine learning algorithms require hundreds of examples to perform similarly. They aim to address a different problem: recognising handwritten characters from image data. In this paper, we position CM2 as a specialized solution optimized for entity extraction (not covered in this paper) and document classification in complex structural domains. By treating the document as a coordinate matrix rather than a linear sequence, CM2 provides the ‘structural intelligence’ necessary for high-stakes industrial processing. Although the issues differ, both works use human-level concept learning by examining how humans solve the problem. Both approaches only use one sample per class and learn from that sample.
@@ 1.3 Advantages
    High Accuracy with Minimal Data (One-Shot Learning): The algorithm achieves human-level concept learning, successfully classifying documents using only a single sample per class. It outperforms traditional machine learning and advanced deep learning models that typically require hundreds or thousands of labeled examples to achieve comparable accuracy. This makes the framework particularly effective for low-data, layout-sensitive tasks where obtaining a large, diverse training corpus is infeasible.
@@
    Geometric and Structural Intelligence: Unlike text-only transformer-based models (like BERT or GPT) that process text as a linear sequence of tokens, CM2 utilizes a coordinate matrix to capture the physical geometry of the document. This makes it significantly more effective for formal documents (bank statements, invoices) where the spatial position of elements is more informative than semantic flow. As a structure-aware document intelligence model, CM2 ensures that even minor spatial deviations are captured, providing a level of layout-sensitivity that standard transformers lack.
@@
    Optimized for CPU-Only Environments: While modern NLP trends often rely on expensive, high-end GPU clusters, CM2 is purpose-built to run on standard consumer-grade CPUs. Using static embeddings, such as GloVe, with $O(1)$ lookup latency rather than complex transformer passes, ensures near-zero inference time on basic hardware.
@@
    Economic Viability: By eliminating the need for expensive GPU infrastructure, high-cost LLM API subscriptions, and the labor-intensive process of large-scale data labeling, CM2 provides a significantly cheaper solution for enterprise-level deployment.
@@
    Generic, Expandable, and Extendable: The coordinate matrix framework is highly versatile. It can be easily adapted to incorporate any structured document templates or extended to include additional feature markers without requiring the model to be “re-trained” on a massive corpus.
@@ 2 Related Work
Once the feature vector is identified, the third step is to apply a machine learning model for classification. Naïve Bayes is one of the probability-based classifiers [kotenko2015]. Decision tree-based approaches have also been used for various purposes, such as blocking inappropriate web content [liu2017]. Support Vector Machines (SVM) are also widely used for document classification [farhoodi2010, lin2006]. There are multiple approaches under artificial/deep neural networks. Hassan and Mahmood [hassan2017] applied Convolution Neural Networks (CNN) for document classification. Unlike most machine learning models, Recurrent Neural Networks (RNNs) consider the sequence of word occurrences; therefore, documents with similar words can yield different outputs because of word order [banerjee2019, kumar2019]. The critical distinction in our methodology is that, whereas most approaches consider text and semantic proximity, we treat both text and coordinates as primary features. This enables CM2 to excel in environments where the physical location of a token is as meaningful as the token itself.